# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

### v2

@@ Abstract
Lyophilization (aka freeze drying) is a typical process in pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines. While extensive efforts have been dedicated to shifting the pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode. This article presents the first mechanistic model for a complete continuous lyophilization process, which comprehensively incorporates and describes key transport phenomena in all three steps of lyophilization, namely freezing, primary drying, and secondary drying. The proposed model considers the state-of-the-art lyophilization technology, in which vials are suspended and move continuously through the process. The validated model can accurately predict the evolution of critical process parameters, including the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire process. Several applications related to model-based process design and optimization of continuous lyophilization are also demonstrated. The final model is made available in MATLAB and Julia as an open-source software package called ContLyo, which can ultimately be leveraged for guiding the design and development of future continuous lyophilization processes.
**footnotetext: Corresponding author. Email: braatz@mit.edu
@@
Keywords— Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
Introduction
@@
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1]. Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3]. Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry. A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and liquid solvent (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [4]. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
@@
Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [5, 6], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [7, 6]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [8] and spin freezing [5] (see a comprehensive review of continuous lyophilization technologies in Ref. [6]). A novel lyophilization technology has been recently proposed by Ref. [7], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
@@
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the key phenomena in lyophilization are underpinned by heat and mass transfer theories, process models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [9, 10, 11, 12, 13], primary drying [14, 15, 16, 17, 18, 9, 19, 20, 21, 4, 22], and secondary drying [14, 15, 16, 17, 18, 23, 4, 24, 25]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref. [26]. A model for the freezing step was also developed for spin freezing [27]. For spray freeze drying, Ref. [28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
@@
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 details the model validation and showcases various applications of the validated model. Finally, Section 5 summarizes the study and briefly discusses possible future work.
Process Description
This work considers lyophilization of unit doses, in which the product is introduced into a vial prior to being lyophilized. This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [6]. Therefore, results and discussion presented in this article are based entirely on lyophilization of unit doses.
Batch lyophilization
The majority of lyophilization processes in the pharmaceutical industry have been operated in a batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig. 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps. In this case, the freezing and drying processes take place in the same space but at different times.
@@
Several drawbacks associated with batch lyophilization are extensively discussed in Refs. [7, 6]. Typical issues of any batch process include batch-to-batch variability, quality control, process downtime, and operational flexibility. Another disadvantage that is specific to the configuration illustrated in Fig. 1A is non-uniformity in heat transfer. For example, vials located on the side and corner of the shelf are typically affected by thermal radiation from the chamber walls, whereas the effect is almost negligible for center vials. Variability in heat transfer conditions could result in different crystal structures and drying times, hence variation in the final product quality. Consequently, process control and optimization are not straightforward. These known issues motivate the development of a continuous lyophilization concept that can be practically implemented in an industrial scale.
Continuous lyophilization
@@
A comprehensive review of continuous lyophilization technologies is given in Ref. [6]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [8, 28]. Both aforementioned approaches have been proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [5].
@@
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref. [7]. This technology employs the suspended-vial configuration as illustrated in Fig. 1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf. Unlike conventional lyophilizers, the freezing and drying processes occur in different locations/chambers of the equipment. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve any complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous.
@@
The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Fig. 1B, the current state-of-the-art continuous lyophilization technology for unit doses. Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in Ref. [7].
Mechanistic Modeling
@@
This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Fig. 1B). The section starts by discussing existing models and modeling strategies in the literature, and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
Review of the existing models and modeling strategies
@@
Mechanistic modeling for conventional/batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [19, 4], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section first provides a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes with the optimal modeling strategies used in this work. The published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
@@
Modeling strategies for the freezing step
@@
Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Fig. 2 highlights notable modeling strategies for this step. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [10, 12].
@@
A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in Ref. [9]. Recently, Ref. [12] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [13]. In Appendix A, we provide a detailed analysis using the relevant dimensionless group to compare between the lumped capacity model and model with thermal gradients and show that the lumped capacity model is sufficiently accurate for lyophilization of unit doses in general.
@@
The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the solid structure and mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [11, 32], how to best model the specific complex interacting phenomena remains an open research question.
Modeling strategies for the primary drying step
@@
Many models for primary drying are available in the literature (see the green column in Fig. 2). The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and hence the temperature gradient in the vertical direction is significant. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 9, 35]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer.
@@
An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref. [19]). In this case, the driving force for mass transfer is dependent on the saturation (equilibrium) pressure at the sublimation interface, which is a function of temperature. Hence, the model should be able to accurately predict the spatial variation of the product temperature, including the interface temperature, and hence modeling the heat transfer in the vertical direction is needed. Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [1], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
@@
There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [15, 16] and in higher dimensions [17]. Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [19, 4]. Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [4]. Consequently, these high-fidelity models are not commonly used.
@@
Note that, in primary drying, the water vapor is removed via both sublimation (frozen region) and desorption (dried region). Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [16, 17, 19, 4].
Modeling strategies for the secondary drying step

### v3

# Mechanistic Modeling of Continuous Lyophilization for Biopharmaceutical Manufacturing

@@ Abstract
Lyophilization (aka freeze drying) is a typical process in (bio)pharmaceutical manufacturing used for improving the stability of various drug products, including its recent applications to mRNA vaccines. While extensive efforts have been dedicated to shifting the (bio)pharmaceutical industry toward continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in batch mode. This article presents the first mechanistic model for a complete continuous lyophilization process, which comprehensively incorporates and describes key transport phenomena in all three steps of lyophilization, namely freezing, primary drying, and secondary drying. The proposed model considers the state-of-the-art lyophilization technology, in which vials are suspended and move continuously through the process. The validated model can accurately predict the evolution of critical process parameters, including the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire process. Several applications related to model-based process design and optimization of continuous lyophilization are also demonstrated. The final model is made available as an open-source software package that can be leveraged for guiding the design and development of future continuous lyophilization processes.
keywords:
@@
Lyophilization, Freeze Drying, Suspended Vials, Continuous Manufacturing, Pharmaceuticals
Prakitr Srisuma Gang Chen Richard D. Braatz*
{affiliations}
P. Srisuma, G. Chen, R. D. Braatz
Massachusetts Institute of Technology, Cambridge, MA 02139, USA
Email: braatz@mit.edu
1 Introduction
Lyophilization, also known as freeze drying, is a low-temperature, low-pressure dehydration process used for improving the stability of various drug products in the biopharmaceutical industry [Fissore2018Review]. By removing the liquid component (usually water) from the product, the final lyophilized product becomes more stable, hence longer shelf life. One of the most recent applications of lyophilization is to provide long-term stability for the mRNA COVID-19 vaccines at room temperature [Muramatsu2022mRNA, Meulewaeter2023mRNA]. This advancement eliminates the need for an ultra-cold supply chain for maintaining the stability of mRNA, which helps facilitate the storage and distribution of mRNA drug product across the world Advancement in lyophilization technology could therefore play a crucial role in the future of mRNA manufacturing and the biopharmaceutical industry in general.
@@ 1 Introduction
A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and water in a vial are cooled such that most of the liquid (free water) is frozen, with the remaining fraction (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals [Fissore2015Review]. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
@@
Although the current trends in the biopharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in batch mode [Meyer2015SpinFreezing, Pisano2019ReviewContLyo], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [Capozzi2019ContLyo_SuspendedVials, Pisano2019ReviewContLyo]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for (bio)pharmaceutical products have been proposed, including spray freeze drying [Ishwarya2015SprayFD] and spin freezing [Meyer2015SpinFreezing] (see a comprehensive review of continuous lyophilization technologies in [Pisano2019ReviewContLyo]). A novel lyophilization technology has been recently proposed by [Capozzi2019ContLyo_SuspendedVials], in which vials are suspended and continuously move along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
@@
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the key phenomena in lyophilization are underpinned by heat and mass transfer theories, process models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [Hottot2006Freezing, Nakagawa2007Freezing, Arsiccio2017CrystalSize, Deck2022FreezingLumped, Deck2024Freezing2D], primary drying [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling, Pikal2005Model, Hottot2006Freezing, Veraldi2008SimBatchModels, Pisano2010control, Chen2015FEMmodel, Fissore2015Review, Bano2020LumpedDrying], and secondary drying [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling, Pikal2005Model, Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by [Bockstal2017ContLyo_ModelPrimary]. A model for the freezing step was also developed for spin freezing [Nuytten2021SpinFreezing]. For spray freeze drying, [Sebastiao2019SprayFD] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
@@
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 discussed the numerical methods required for solving and simulating the model equations efficiently. Section 5 details the model validation and showcases various applications of the validated model. Finally, Section 6 summarizes the study and briefly discusses possible future work.
2 Process Description
This work considers lyophilization of unit doses, in which the product is introduced into a vial prior to being lyophilized. This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [Pisano2019ReviewContLyo]. Therefore, results and discussion presented in this article are based entirely on lyophilization of unit doses.
2.1 Batch lyophilization
The majority of lyophilization processes in the (bio)pharmaceutical industry have been operated in batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Figure 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps. In this case, the freezing and drying processes take place in the same space but at different times.
@@ 2.1 Batch lyophilization
Several drawbacks associated with batch lyophilization are extensively discussed in [Capozzi2019ContLyo_SuspendedVials, Pisano2019ReviewContLyo]. Typical issues of any batch process include batch-to-batch variability, quality control, process downtime, and operational flexibility. Another disadvantage that is specific to the configuration illustrated in Figure 1A is non-uniformity in heat transfer. For example, vials located on the side and corner of the shelf are typically affected by thermal radiation from the chamber walls, whereas the effect is almost negligible for center vials. Variability in heat transfer conditions could result in different crystal structures and drying times, hence variation in the final product quality. Consequently, process control and optimization are not straightforward. These known issues motivate the development of a continuous lyophilization concept that can be practically implemented in an industrial scale.
2.2 Continuous lyophilization
@@
A comprehensive review of continuous lyophilization technologies is given in [Pisano2019ReviewContLyo]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [Gottfried1973ContLyoPatent] or a revolving plate with holes [Bruttini1993ContLyoPatent]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with a high surface-to-volume ratio, greatly improving heat and mass transfer in the process [Ishwarya2015SprayFD, Sebastiao2019SprayFD]. Both aforementioned approaches have been proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [Meyer2015SpinFreezing].
@@
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by [Capozzi2019ContLyo_SuspendedVials]. This technology employs the suspended-vial configuration as illustrated in Figure 1B, where all vials are suspended and continuously move through the freezing and drying chambers, without any contact between the vials and shelf. Unlike conventional lyophilizers, the freezing and drying processes occur in different locations/chambers of the equipment. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve any complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in (bio)pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous.
@@
The mechanistic model developed in this work is based on this suspended-vial configuration as illustrated in Figure 1B, the current state-of-the-art continuous lyophilization technology for unit doses. Detailed modeling strategies and model equations are discussed in the next section (Section 3), while information and discussion related to the equipment design can be found in [Capozzi2019ContLyo_SuspendedVials].
3 Mechanistic Modeling
@@
This section details the development of a mechanistic model for continuous lyophilization of suspended vials (Figure 1B). The section starts by discussing existing models and modeling strategies in the literature, and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
3.1 Review of the existing models and modeling strategies
@@
Mechanistic modeling for conventional/batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. While some past reviews briefly discuss the literature on mechanistic modeling of lyophilization [Veraldi2008SimBatchModels, Fissore2015Review], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section first provides a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes with the optimal modeling strategies used in this work. The published models for continuous lyophilization mentioned in Section 1 are based on the spin and spray freeze drying, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
@@
3.1.1 Modeling strategies for the freezing step
@@
Modeling the freezing step is most complicated among the three steps of lyophilization; the blue column in Figure 2 highlights notable modeling strategies for this step. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [Mills1995HeatTransfer]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails supercooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point [Nakagawa2007Freezing, Deck2022FreezingLumped].
@@
A more common but more complicated strategy is to consider the supercooling period and incorporate the heat transfer associated with ice nucleation as done in [Hottot2006Freezing]. Recently, [Deck2022FreezingLumped] introduced a state-of-the-art freezing model that incorporates the stochastic nature of ice nucleation [Colucci2020CrystalSize] and considers heat transfer among multiple vials in batch lyophilization using a lumped capacity model. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [Deck2024Freezing2D]. In Appendix A, we provide a detailed analysis using the relevant dimensionless group to compare between the lumped capacity model and model with thermal gradients and show that the lumped capacity model is sufficiently accurate for lyophilization of unit doses in general.
@@
The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the solid structure and mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [Arsiccio2017CrystalSize, Colucci2020CrystalSize], how to best model the specific complex interacting phenomena remains an open research question.
3.1.2 Modeling strategies for the primary drying step
@@
Many models for primary drying are available in the literature (see the green column in Figure 2). The simplest strategy is to use a lumped capacity model [Bano2020LumpedDrying], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and hence the temperature gradient in the vertical direction is significant. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [Dyer1968HeatTransferLimit, Jafar2003HeatTransferLimit, Hottot2006Freezing, Srisuma2023Analytical]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [Srisuma2023Analytical], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer.
@@
An approach to incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front/interface between the frozen and dried regions, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of [Veraldi2008SimBatchModels]). In this case, the driving force for mass transfer is dependent on the saturation (equilibrium) pressure at the sublimation interface, which is a function of temperature. Hence, the model should be able to accurately predict the spatial variation of the product temperature, including the interface temperature, and hence modeling the heat transfer in the vertical direction is needed. Moreover, one of the key considerations during the drying steps is to ensure that the product temperature does not exceed the upper limit (e.g., collapse temperature and glass transition temperature) [Fissore2018Review], and thus spatially distributed temperature data are valuable for accurately determining the maximum temperature in the product.
@@
There are models that consider detailed heat and mass transfer by including full energy and continuity equations, both in 1D [Liapis1994Original, Sadikoglu1997Modeling] and in higher dimensions [Sheehan1998Modeling]. Implementing such high-fidelity models for practical applications, e.g., state estimation and control, is not easy due to their high complexity, computation time, and number of parameters [Veraldi2008SimBatchModels, Fissore2015Review]. Besides, the accuracy of the complex models for primary drying is not significantly different from 1D or simplified models in most cases [Fissore2015Review]. Consequently, these high-fidelity models are not commonly used.
@@
Note that, in primary drying, the water vapor is removed via both sublimation (frozen region) and desorption (dried region). Nevertheless, it is widely acknowledged that the amount of water vapor removed via sublimation is much higher than that of desorption, and so the effect of desorption can be omitted [Sadikoglu1997Modeling, Sheehan1998Modeling, Veraldi2008SimBatchModels, Fissore2015Review].
3.1.3 Modeling strategies for the secondary drying step